OpenStudy (anonymous):

what is the to 7/8+2/3+5/6=reduced

OpenStudy (anonymous):

7/3

OpenStudy (anonymous):

You must find a common denomanator first which is 24

OpenStudy (anonymous):

ok 24 but the answer cant be 7/3 reduced

OpenStudy (hoblos):

19/8

OpenStudy (anonymous):

how do u get 19/8

OpenStudy (hoblos):

it is 21/24 +16/24 +20/24 =57/24 =19/8

OpenStudy (anonymous):

7/8=21/24 2/3=16/24 5/6=20/24

OpenStudy (anonymous):

Oh ya right when I added i got 56/24

OpenStudy (anonymous):

I am wrong Listen to hobbles

OpenStudy (anonymous):

it still cant be 19/8 simplified

OpenStudy (hoblos):

19/8 is the simplified form of your expression

OpenStudy (anonymous):

It can be written as a mixed number

OpenStudy (anonymous):

no it isnt 19 over 8 is to big

OpenStudy (anonymous):

yea i mean the mixed number

OpenStudy (anonymous):

2 3/8

OpenStudy (anonymous):

thaniks